# Inspection-Guided Randomization: A Flexible and Transparent Restricted Randomization Framework for Better Experimental Design

### v1

@@ Abstract
Randomized experiments are considered the gold standard for estimating causal effects. However, out of the set of possible randomized assignments, some may be likely to produce poor effect estimates and misleading conclusions. Restricted randomization is an experimental design strategy that filters out undesirable treatment assignments, but its application has primarily been limited to ensuring covariate balance in two-arm studies where the target estimand is the average treatment effect. Other experimental settings with different design desiderata and target effect estimands could also stand to benefit from a restricted randomization approach. We introduce Inspection-Guided Randomization (IGR), a transparent and flexible framework for restricted randomization that filters out undesirable treatment assignments by inspecting assignments against analyst-specified, domain-informed design desiderata. In IGR, the acceptable treatment assignments are locked in ex ante and pre-registered in the trial protocol, thus safeguarding against p-hacking and promoting reproducibility. Through illustrative simulation studies motivated by education and behavioral health interventions, we demonstrate how IGR can be used to improve effect estimates compared to benchmark designs in group formation experiments and experiments with interference.
@@ 1 Introduction
This work is organized as follows. In Section 2, we review prior work in restricted randomization and experimental design, then highlight the contributions we make through IGR. In Section 3, we present each step of the IGR framework in detail. In Section 4, we use illustrative vignettes with simulations to show how IGR can be applied to the design of group formation experiments and experiments with interference. In Section 5, we summarize the benefits of using IGR and propose directions for future work.
@@ 2 Related Work
Instead of listing all possible allocations at once, re-randomization repeatedly randomizes until producing the first allocation that meets pre-specified criteria (Rubin,, 2008; Cox,, 2009; Morgan and Rubin,, 2012). Re-randomization has been used beyond two-arm study designs in experiments with factorial designs (Branson et al.,, 2016) and in experiments on networks where units have correlated potential outcomes (Basse and Airoldi,, 2018). Although re-randomization can save computing power, it opens the door to p-hacking because the set of criteria-meeting allocations used to conduct a randomization test are sampled after study outcomes have been observed. Additionally, re-randomization assumes that the analyst’s first choice of restriction criteria produces a satisfactory design, when in reality the criteria may permit allocations that have poor properties for estimation and inference.
@@
IGR unifies and builds upon ideas across several of the works above. The steps of IGR closely resemble those of covariate-constrained randomization (Moulton,, 2004) but, unlike covariate-constrained randomization, IGR recognizes that it can be beneficial to inspect allocations on the basis of multiple properties, particularly properties other than balance. IGR is also related to model-assisted design (Basse and Airoldi,, 2018), but whereas model-assisted design requires deriving a closed-form expression for mean-squared error, IGR can use domain-informed, approximate metrics when the expression for mean-squared error is unknown or difficult to derive. Furthermore, in line with efforts to make studies more reproducible, IGR includes a pre-registration step, which is not explicitly present in any existing restricted randomization framework. Finally, IGR places an emphasis on deliberately interrogating the quality of the restricted design and on making subsequent adaptations to the restriction as necessary. This adaptation step is either under-emphasized or absent from prior approaches to restricted randomization. In practice, we suspect that two common forms of adaptation will be: (i) fixing inspection metrics that do not correctly target the design desiderata (e.g., a balance metric that does not capture the kind of imbalance concerning to the study stakeholders) and (ii) adjusting the relative emphases placed on different design desiderata to better navigate trade-offs (e.g., in trials with interference, homophily can create tension between achieving covariate balance and minimizing interference, and which of the two desiderata is prioritized may need to be adjusted). We explore the latter form of adaptation in Vignette 2 (Section 4.2) and discuss further in Section 5.
3 Inspection-Guided Randomization
@@ 3.1 Setup and notation
We consider a setting with $N$ participants in a $k$-armed study, where the participants are indexed by $i=1,...,N$. Let $\mathbf{X}$ be the $N$-by-$p$ dimensional matrix representing the participants’ baseline covariates, where we denote the covariate vector for participant $i$ as $X_{i\cdot}$ and the vector of values for covariate $j$ as $X_{\cdot j}$. For a given participant $i$, we use the notation $Z_{i}$ to denote the participant’s treatment assignment, where $Z_{i}\in\{0,...,k-1\}$. Let the $N$-dimensional vector of treatments be denoted $\mathbf{Z}=(Z_{1},Z_{2},\dots,Z_{N})$. Since treatment is assigned at random, $Z_{i}$ is a random variable, and it follows that $\mathbf{Z}$ is a random vector. We denote a particular realization of a treatment assignment for individual $i$ as lowercase $z_{i}$ and likewise denote the intervention assignment vector lowercase $\mathbf{z}$. The probability that random vector $\mathbf{Z}$ takes on a particular value $\mathbf{z}$ is written $P(\mathbf{Z}=\mathbf{z})=p_{\mathbf{z}}$, and we refer to $P(\mathbf{Z})$ as the assignment mechanism. We call a particular realization of the intervention assignment vector a treatment allocation. Letting $\mathcal{Z}$ denote the collection of all $k^{N}$ possible values of $\mathbf{z}$, we have that $\sum_{\mathbf{z}\in\mathcal{Z}}p_{\mathbf{z}}=1$.
3.2 Designing experiments with IGR: A step-by-step framework
For clarity, we organize the IGR design framework into the following steps: 1. Specification, 2. Enumeration, 3. Restriction, 4. Evaluation & Adaptation, 5. Pre-registration, and 6. Randomization. We elaborate on each of these steps below.
1. Specification
@@
The study analyst first works with the domain experts on the research team to specify the goals of the study in terms of the causal question of interest and the target effect estimand. Based on these goals, the team compiles desiderata for the study such that if the deployed randomization allocation were guaranteed to satisfy all desiderata, then the data collected would likely be conducive to credible causal inference.
@@
The analyst then constructs a set of inspection metrics that are used to assess candidate allocations, where each metric measures the extent to which the allocation satisfies a particular desideratum (e.g., standardized mean difference as a metric to measure balance). We assume that these metrics are defined such that they map each allocation and observed covariates to a scalar value. We denote a collection of $L$ metrics as $m_{1},m_{2},\dots,m_{L}$ and denote $s_{1},s_{2},\dots s_{L}$ as the scalar outputs when these metrics are applied to inspect a particular allocation $\mathbf{z}$, where $s_{l}=m_{l}(\mathbf{z},\mathbf{X})\in\mathbb{R},\ l=1,2,\dots,L$.
@@
The analyst also decides on an aggregator function, $\Theta$, that aggregates $s_{1},s_{2},\dots s_{L}$ into a single score that summarizes an allocation’s desirability. We write $\Theta(s_{1},s_{2},\dots,s_{L})=s$, where $s\in\mathbb{R}$. In certain settings, the desiderata may be conflicting, where scoring better for some desiderata necessarily entails worse scores on other desiderata. The aggregator can be used to navigate such desiderata trade-offs, for instance by incorporating a weighting scheme that prioritizes certain desiderata over others. Let $f$ be the function that maps an allocation $\mathbf{z}$ to its score by combining the outputs of each of the individual inspection metric via the aggregation function $\Theta$, that is $f(\mathbf{z},\mathbf{X})=\Theta(m_{1}(\mathbf{z},\mathbf{X}),m_{2}(\mathbf{z},\mathbf{X}),\dots,m_{L}(\mathbf{z},\mathbf{X})).$ We refer to $f$ as the fitness function.
@@
2. Enumeration
The analyst enumerates a pool of unique candidate treatment allocations, $\mathcal{Z}_{pool}=\{\mathbf{z_{1}},\mathbf{z_{2}},\dots\mathbf{z_{M}}\}$. For a $k$-armed study, there are $k^{N}$ total possible allocations. When $N$ is large, it may be necessary to set $M<<k^{N}$. The enumeration of $\mathcal{Z}_{pool}$ can be done by sampling from any choice of assignment mechanism $P$, including Bernoulli randomization and complete randomization. In some cases, sampling from an assignment mechanism that is already restricted, such as a block randomization, can increase the efficiency of the enumeration step, since the mechanism inherently rules out some of the undesirable allocations. Similarly, it is also possible to use enumeration strategies that incorporate optimization routines and more efficiently search the sample space for allocations that have good fitness function scores (see Section 3.3).
3. Restriction
Applying the fitness function $f$ and restriction rule $r$ to the enumerated pool of candidate allocations yields the following assignment mechanism:
@@ 3. Restriction
At this step, the assignment mechanism $P_{IGR}$ is proposed but not yet locked in.
4. Evaluation & Adaptation
Before proceeding and locking in the restricted assignment mechanism, the study analyst and domain experts pause to evaluate the fitness function $f$ and restriction rule $r$, making adaptations and improvements to them if necessary. Even if $f$ and $r$ were carefully identified in the Specification, they still may be flawed in unforeseen ways and could yield poor study designs. We highlight three dimensions along which $f$ and $r$ should be evaluated: discriminatory power, desiderata trade-offs, and over-restriction.

### v2

@@ Abstract
Randomized experiments are considered the gold standard for estimating causal effects. However, out of the set of possible randomized assignments, some may be likely to produce poor effect estimates and misleading conclusions. Restricted randomization is an experimental design strategy that filters out undesirable treatment assignments, but its application has primarily been limited to ensuring covariate balance in two-arm studies where the target estimand is the average treatment effect. Other experimental settings with different design desiderata and target effect estimands could also stand to benefit from a restricted randomization approach. We introduce Inspection-Guided Randomization (IGR), a transparent and flexible framework for restricted randomization that filters out undesirable treatment assignments by inspecting assignments against analyst-specified, domain-informed design desiderata. In IGR, the acceptable treatment assignments are locked in ex ante and pre-registered in the trial protocol, thus safeguarding against p-hacking and promoting reproducibility. Through illustrative simulation studies motivated by behavioral health and education interventions, we demonstrate how IGR can be used to improve effect estimates compared to benchmark designs in experiments with interference and in group formation experiments.
@@ 1 Introduction
This work is organized as follows. In Section 2, we review prior work in restricted randomization and experimental design, then highlight the contributions we make through IGR. In Section 3, we introduce notation and an overview of the IGR framework. In Section 4, we walk through each step of IGR in detail and use a simulation to illustrate how it can be applied to the design of an experiment with interference. In Sections 5-6, we discuss analysis and inference for IGR-designed experiments and show how IGR can improve effect estimates. We conclude in Section 7 with a summary of the benefits of using IGR and directions for future work.
@@ 2 Related Work
Instead of listing all possible allocations at once, re-randomization repeatedly randomizes until producing the first allocation that meets pre-specified criteria (Rubin,, 2008; Cox,, 2009; Morgan and Rubin,, 2012). Re-randomization has been used beyond two-arm study designs in experiments with factorial designs (Branson et al.,, 2016) and in experiments on networks where units have correlated potential outcomes (Basse and Airoldi,, 2018). Although re-randomization can save computing power, it opens the door to p-hacking because the set of criteria-meeting allocations used to conduct a randomization test are sampled after study outcomes have been observed (see further discussion in Appendix G on the risk of Type I error rate inflation when the accepted allocations in the analysis phase are different from those in the design phase). Additionally, re-randomization assumes that the analyst’s first choice of restriction criteria produces a satisfactory design, when in reality the criteria may permit allocations that have poor properties for estimation and inference.
@@
IGR unifies and builds upon ideas across several of the works above. Tightening, re-randomization, and covariate-constrained randomization were developed to target balance criteria only, and subsequent literature that applies or expands these randomization strategies also only consider balance (Appendix Table A.1). While IGR resembles covariate-constrained randomization, we make an important extension by recognizing that it can be beneficial to inspect allocations on the basis of multiple properties, particularly properties other than balance. IGR is also related to model-assisted design (Basse and Airoldi,, 2018), but whereas model-assisted design requires deriving a closed-form expression for mean-squared error, IGR can use domain-informed, approximate metrics when the expression for mean-squared error is unknown or difficult to derive.
Additionally, most prior restricted randomization strategies, e.g. re-randomization, use a naive sampling approach to select an allocation. In contrast, in IGR, we can efficiently “search” through a large candidate allocation space (e.g., with genetic algorithms) to enumerate a set of candidate allocations that score well according to specified metrics. Speaking more generally, in re-randomization, the process of searching for an allocation and the process of randomizing are conflated in the act of sampling. IGR, on the other hand, separates the enumeration step, which can be done without any randomization, from the randomization step, and thus draws a clear distinction between searching and randomizing. Some implementations of constrained randomization use a similar approach to IGR (Moulton,, 2004; de Hoop et al.,, 2012; Li et al.,, 2016; Ciolino et al.,, 2019; Watson et al.,, 2021; Zhou et al.,, 2021), where a candidate set of allocations are first enumerated, and the final allocation is chosen randomly from among criterion-meeting allocations within the candidate set. However, these implementations use a random sampling procedure to produce the initial candidate set; they do not highlight that enumerating the candidate set can involve a directed search and does not necessarily need to be random. The randomness used for inference occurs later, in the actual assignment.
In line with efforts to make studies more reproducible, IGR includes and enforces a step where accepted allocations are pre-registered. As long as the study analyst performs randomization inference over the pre-registered set of accepted allocations, then getting the nominal Type I error rate is guaranteed. Pre-registration is not explicitly present in any existing restricted randomization framework, so study analysts are not held accountable to conducting correct analyses. We show, for instance, that in re-randomization, we can easily inflate the Type I error rate by using a more restrictive balance criterion in the analysis phase than in the design phase.
@@
Finally, IGR places an emphasis on deliberately interrogating the quality of the restricted design and on making subsequent adaptations to the restriction as necessary. This adaptation step is either under-emphasized or absent from prior approaches to restricted randomization. In practice, we suspect that two common forms of adaptation will be: (i) fixing inspection metrics that do not correctly target the design desiderata (e.g., a balance metric that does not capture the kind of imbalance concerning to the study stakeholders) and (ii) adjusting the relative emphases placed on different design desiderata to better navigate trade-offs (e.g., in trials with interference, homophily can create tension between achieving covariate balance and minimizing interference, and which of the two desiderata is prioritized may need to be adjusted).
Several prior real-world studies have been designed with restricted randomization techniques related to IGR that included some, but not all, of the elements of the IGR framework (Murnane et al.,, 2023; Cho et al.,, 2021, 2022; Kizilcec et al.,, 2022; Zahrt et al.,, 2023). These studies demonstrate the need for a formal framework like IGR that unifies design principles, and they also highlight the potential for IGR to be applicable and useful in practice. In this work, we present simulated vignettes of an experiment with interference and of a group formation experiment (Section 4 and Appendix E, respectively) to explain specific features of the IGR framework and to highlight how IGR can be used for diverse experimental settings, including and beyond those where related restricted randomization techniques have been used in the past.
3 An Overview of Inspection-Guided Randomization
@@ 3.1 Setup and notation
We consider a setting with $N$ participants in a $k$-armed study, where the participants are indexed by i=1,,Ni=1,...,N. Let $\mathbf{X}$ be the NN-by-$p$ dimensional matrix representing the participants’ baseline covariates. For a given participant $i$, we use the notation $Z_{i}$ to denote the participant’s treatment assignment, where $Z_{i}\in\{0,...,k-1\}$. Let the $N$-dimensional vector of treatments be denoted $\mathbf{Z}=(Z_{1},Z_{2},\dots,Z_{N})$. Since treatment is assigned at random, $Z_{i}$ is a random variable, and it follows that $\mathbf{Z}$ is a random vector. We denote a particular realization of a treatment assignment for individual $i$ as lowercase $z_{i}$ and likewise denote the intervention assignment vector lowercase $\mathbf{z}$. The probability that random vector $\mathbf{Z}$ takes on a particular value $\mathbf{z}$ is written $P(\mathbf{Z}=\mathbf{z})=p_{\mathbf{z}}$, and we refer to $P(\mathbf{Z})$ as the assignment mechanism. We call a particular realization of the intervention assignment vector a treatment allocation. Letting $\mathcal{Z}$ denote the collection of all $k^{N}$ possible values of $\mathbf{z}$, we have that $\sum_{\mathbf{z}\in\mathcal{Z}}p_{\mathbf{z}}=1$.
3.2 Outline of the IGR framework
For clarity and ease of use in practice, we organize the IGR design framework into distinct steps. We describe each in brief below, then elaborate further in the next section:
  1. 1.
    Specification. The analyst specifies the target effect estimand and the fitness function used to measure the desirability of each candidate treatment allocation.
  2. 2.
    Enumeration. The analyst enumerates a large pool of candidate allocations using a simple assignment mechanism.
  3. 3.
    Restriction. The analyst scores each candidate allocation with the specified fitness function and filters out those that are low-scoring, according to a restriction rule.
  4. 4.
    Evaluation & Adaptation. The analyst evaluates their choice of fitness function and restriction rule, then makes adaptations if needed to better achieve design desiderata.
  5. 5.
    Pre-registration. The analyst pre-registers the IGR assignment mechanism.
  6. 6.
    Randomization. The analyst samples one allocation to deploy for the experiment.
4 IGR in Detail, with Application to a Simulated Experiment With Interference
In this section, we provide a full description and rationale for the steps of the IGR framework. To make the steps more concrete, we also illustrate how to apply them to a simulated experiment with interference. We begin this section by describing the domain and set-up for the simulated experiment, then walk through the steps of IGR in detail, explaining how each can be applied in the context of the simulated experiment. Note that IGR can be used in many experimental settings, not just those with interference. In Appendix E, for example, we also highlight how IGR can be applied to a group formation experiment. All simulation code is available at https://github.com/wangmagg/inspection-guided-randomization.
4.1 Simulated Experiment with Interference
The challenge of interference
Interference occurs when the treatment assigned to one observational unit affects the outcome of a different unit. Interference is often present with behavioral interventions, where information can diffuse or spill over along social ties (Valente,, 2012; Cai et al.,, 2015; Kim et al.,, 2015). When interference is present, a common target treatment effect estimand is the global average treatment effect, or a comparison of the average outcome if everyone were to receive treatment versus if everyone were to receive control (Hudgens and Halloran,, 2008). Failing to properly account for interference can make estimates of the global average treatment effect both biased and imprecise (Eckles et al.,, 2017). This in turn can result in misleading conclusions (e.g., false nulls) and misinformed recommendations for policy-making (e.g., failure to implement an effective intervention).
A common approach to controlling interference is to use cluster randomization, where randomization is done on the cluster level such that all individuals in the same cluster receive the same treatment (Hayes and Moulton,, 2017). Given enough separation between clusters, there is likely no spillover between different treatment arms, and comparing treated to control clusters can give an unbiased estimate of the global average treatment effect. However, it may sometimes be difficult to identify clusters that are sufficiently separated to avoid any inter-cluster interference. In other cases, it may be possible to identify a few large clusters that are well-separated, such as villages that are far apart geographically. However, randomizing only a few clusters leads to low power, especially if there is strong intra-cluster correlation on prognostic covariates. Often, for clusters that form organically, individuals within a cluster are more similar than those in different clusters, so intra-cluster correlation is expected.
Within the IGR design framework, we can be more flexible with the way that we control interference. Leveraging prior knowledge, which may not necessarily be complete or fully correct, we can use an interference-targeting constraint to control the amount of anticipated interference in a more precise way than with cluster randomization. Even if we do not have perfect knowledge of the interference structure (e.g., which individuals interact to produce spillover), we can still develop “coarse” metrics that measure interference.
Simulation set-up
4.2 Step-by-Step Application of IGR
We now go through IGR step-by-step in detail. After each step is described, we show how it can be applied to designing the simulated experiment.
4.2.1 Specification
@@
In the Specification step, the study analyst first works with the domain experts on the research team to specify the goals of the study in terms of the causal question of interest and the target effect estimand. Based on these goals, the team compiles desiderata for the study such that if the deployed randomization allocation were guaranteed to satisfy all desiderata, then the data collected would likely be conducive to credible causal inference.
@@
The analyst then constructs a set of inspection metrics that are used to assess candidate allocations, where each metric measures the extent to which the allocation satisfies a particular desideratum (e.g., standardized mean difference as a metric to measure balance). In Table 1, we list three broad categories of common design desiderata and examples of inspection metrics that might be used within each category.
Desideratum Category Example Desideratum Example Inspection Metric(s)
Estimation properties Low variance Standardized mean difference, Mahalanobis distance
Low bias in the presence of network or spatial spillover Connectivity in a graph network, geographic proximity between treated/control units
Study logistics Ease of delivering a treatment for an infectious disease Travel time or distance to reach all units assigned to the treatment arm
Conditions for testing scientific theory Placing individuals in certain social contexts to study effect on academic achievement Compositional characteristics of the individuals assigned to a given classroom (e.g. gender ratio)
Table 1: Common desiderata categories, specific examples of desiderata within each category, and example inspection metrics that can be used to measure the extent to which an allocation satisfies a particular desideratum.
@@
The analyst also decides on an aggregator function that combines the separate metric scores and summarizes an allocation’s desirability. In certain settings, the desiderata may be conflicting, where scoring better for some desiderata necessarily entails worse scores on other desiderata. The aggregator can be used to navigate such desiderata trade-offs, for instance by incorporating a weighting scheme that prioritizes certain desiderata over others. Let $f$ be the fitness function that maps an allocation $\mathbf{z}$ to its score by combining the outputs of each of the individual inspection metrics via the aggregation function.
@@
{leftrule}
Specification in a Simulated Experiment
We specify the target effect estimand to be the global average treatment effect, $\tau_{GATE}=\frac{1}{N}\sum_{i=1}^{N}Y_{i}(\mathbf{1})-Y_{i}(\mathbf{0})$ where $\mathbf{1}$ is the all-ones vector (everyone is treated) and $\mathbf{0}$ is the all-zeros vector (everyone is control). Our design goal is to assign treatments such that those in the treatment (control) arm experience a universe that is as close as possible to an all-treated (all-control) universe while also balancing covariates across arms. Thus, we specify two inspection metrics, one to measure interference and one to measure balance.
[…]
where $z_{i},z_{j}\in\{0,1\}$ are the treatment assignments for units $i$ and $j$, respectively and $\mathbf{l}_{i},\mathbf{l}_{j}\in\mathbb{R}^{2}$ are coordinates for $i$ and $j$. While we use Euclidean distance here, other distance functions can also be used. It is also possible to expand upon a simple distance function by incorporating known information about ease of travel, such as the presence or absence of interconnecting roads. For balance, we use Mahalanobis distance,
$m^{(\texttt{Mahalanobis})}(\mathbf{z},\mathbf{g},\mathbf{X})$ $=\left(\bar{\mathbf{X}}^{(1)}-\bar{\mathbf{X}}^{(0)}\right)\mathbf{S}^{-1}\left(\bar{\mathbf{X}}^{(1)}-\bar{\mathbf{X}}^{(0)}\right)^{T}$ (2)
where $\bar{\mathbf{X}}^{(1)}$ and $\bar{\mathbf{X}}^{(0)}$ are the mean covariate vectors in the treated and control arms, respectively, and $\mathbf{S}$ is the sample covariance matrix. Combining the interference and balance metrics, we define the fitness function
$\begin{split}f^{(\texttt{I+M})}(\mathbf{z},\mathbf{X})&=0.5\cdot m^{(\texttt{InvMinEuclideanDist})}(\mathbf{z},\mathbf{X})+0.5\cdot m^{(\texttt{Mahalanobis})}\cdot(\mathbf{z},\mathbf{X})\end{split}$ (3)
To put the metrics on the same scale, we standardize the values across all enumerated candidate allocations (see the Enumeration step below) before combining the metrics in the fitness function.
4.2.2 Enumeration
In the Enumeration step, the analyst enumerates a pool of unique candidate allocations, $\mathcal{Z}_{pool}=\{\mathbf{z_{1}},\mathbf{z_{2}},\dots\mathbf{z_{M}}\}$. For a $k$-armed study, there are $k^{N}$ total possible allocations. When $N$ is large, it may be necessary to set $M<<k^{N}$. The enumeration of $\mathcal{Z}_{pool}$ can be done by sampling from any choice of assignment mechanism $P$, including Bernoulli randomization and complete randomization. In some cases, sampling from an assignment mechanism that is already restricted, such as a block randomization, can increase the efficiency of the enumeration step, since the mechanism inherently rules out some of the undesirable allocations (see Appendix E for an example).
Since $M<<k^{N}$, it is possible that many desirable, high-scoring allocations are overlooked and not enumerated as part of $\mathcal{Z}_{pool}$. Instead of using simple enumeration strategies, we may instead want to efficiently search through the space of all $k^{N}$ possible allocations to construct a more optimal $\mathcal{Z}_{pool}^{(*)}$. One approach to construct $\mathcal{Z}^{(*)}_{pool}$ is to use genetic algorithms (Mitchell,, 1998). The initial $\mathcal{Z}_{pool}$ is viewed as a population of treatment allocations. These allocations “mate” in pairs to produce a new generation of allocations, where novel allocations are introduced through “mutation” (single entries of an allocation are changed) and “crossover” (segments of two parent allocations are swapped). The individuals in this generation are scored using the specified fitness function $f$, and those with the best scores survive and breed to produce the next generation. We can repeat this process over $T$ generations to obtain $\mathcal{Z}_{pool}^{(*)}$.
{leftrule}
Enumeration in a Simulated Experiment
We first sample candidate allocations from a complete randomization assignment mechanism that assigns an equal number of schools to the treatment and control arms. We enumerate a total of $M=100,000$ allocations in the initial candidate pool, which is in line with prior work (Li et al.,, 2016; Watson et al.,, 2021; Nordin and Schultzberg,, 2022). We then apply 3 generations of mutations and crossover to the initially enumerated pool to search for a pool with better fitness scores.
4.2.3 Restriction
The analyst selects a restriction rule $r$, then applies the fitness function and restriction rule to the enumerated pool of candidate allocations to produce the following assignment mechanism:
@@ 4.2.3 Restriction
We refer to allocations that have a non-zero probability of being sampled as “accepted” allocations. At this step, the assignment mechanism $P_{IGR}$ is proposed but not yet locked in.
{leftrule}
Restriction in a Simulated Experiment
We use the restriction rule $r(s)=0\text{ if }\ s\geq s^{*},r(s)=1\text{ if }s<s^{*}$, where $s^{*}$ is the $0.5^{\text{th}}$-percentile of scores for the enumerated candidate allocations. This restriction rule filters out all but the top $0.5^{\text{th}}$-percentile of the $100,000$ enumerated allocations, resulting in $m=500$ accepted allocations.
4.2.4 Evaluation & Adaptation
Before proceeding and locking in the restricted assignment mechanism, the study analyst and domain experts pause to evaluate the fitness function $f$ and restriction rule $r$, making adaptations and improvements to them if necessary. Even if $f$ and $r$ were carefully identified in Specification, they still may be flawed in unforeseen ways and could yield poor study designs. We highlight three dimensions along which $f$ and $r$ should be evaluated: discriminatory power, desiderata trade-offs, and over-restriction.